OpenStudy (anonymous):

How do I find the slope of the line for 3x+5y-15=0?

OpenStudy (anonymous):

5y=-3x+15 y=(-3/5)x+3 which is in the form y=mx+c so slope=-3/5

OpenStudy (anonymous):

Solve the equation for y. The slope will be the coefficient (number) in front of the x variable,.

OpenStudy (anonymous):

thank you